OpenStudy (anonymous):

HI AGAIN Candice deposited $3,587.39 into a savings account with an interest rate of 4.2% compounded twice a year. About how long will it take for the account to be worth $5,000?

OpenStudy (anonymous):

@jim_thompson5910 hi

jimthompson5910 (jim_thompson5910):

now you use a similar formula

jimthompson5910 (jim_thompson5910):

A = P(1+r/n)^(n*t) in this case A = 5000 P = 3587.39 r = 0.042 n = 2 (compounding 2 times a year) t = unknown plug all that in and solve for t

OpenStudy (anonymous):

i hate solving for crap algebraically, it makes my head bleed...

jimthompson5910 (jim_thompson5910):

well this is an algebra class (sounds like it?) so that's to be expected the solving part, not head bleeding part

OpenStudy (anonymous):

i know lol

OpenStudy (anonymous):

im trying

jimthompson5910 (jim_thompson5910):

ok tell me what you get

OpenStudy (anonymous):

i got this far\[1412.61=(1.021)^{2timest}\]

jimthompson5910 (jim_thompson5910):

how are you getting 1412.61?

OpenStudy (anonymous):

subtracting 3587.39 from 5000

OpenStudy (anonymous):

its probably soooooo wrong

jimthompson5910 (jim_thompson5910):

oh...you don't subtract you're actually dividing

OpenStudy (anonymous):

thats why im here

jimthompson5910 (jim_thompson5910):

close though

jimthompson5910 (jim_thompson5910):

yeah that's why it's best to show any/all work you can

OpenStudy (anonymous):

ummmmm... i dont know what to do

OpenStudy (anonymous):

so i divide what?

jimthompson5910 (jim_thompson5910):

5000 = 3587.39(1.021)^(2*t) divide both sides by 3587.39

OpenStudy (anonymous):

i dont have a texas instrument calculator, i have a phone

jimthompson5910 (jim_thompson5910):

just do 5000/3587.39 and tell me what you get

OpenStudy (anonymous):

i cant do this on my phone

OpenStudy (anonymous):

i can do that

jimthompson5910 (jim_thompson5910):

btw you shouldn't be doing this on a phone lol you need a real calculator

OpenStudy (anonymous):

i know, my friend had to borrow mine haha

jimthompson5910 (jim_thompson5910):

google or wolfram alpha are good choices example if you typed in "5000/3587.39" without quotes into google, you would get https://www.google.com/search?hl=&q=5000%2F3587.39&sourceid=navclient-ff&rlz=1B3GGLL_enUS420US420&ie=UTF-8

jimthompson5910 (jim_thompson5910):

or you can use calculators for the computer (like graphcalc)

OpenStudy (anonymous):

and i got 1.39 for 5000/3587.39

OpenStudy (anonymous):

cool, didnt know that

jimthompson5910 (jim_thompson5910):

yeah roughly 1.39377, good

jimthompson5910 (jim_thompson5910):

so 1.39377 = (1.021)^(2*t)

OpenStudy (anonymous):

okay

jimthompson5910 (jim_thompson5910):

how do you isolate t?

OpenStudy (anonymous):

dang man, i forget, i havent taken algebra since sophomore year, im a senior

jimthompson5910 (jim_thompson5910):

so it's been 2 years

OpenStudy (anonymous):

i probably sound really stupid haha

jimthompson5910 (jim_thompson5910):

that's a long time

jimthompson5910 (jim_thompson5910):

nah not really

OpenStudy (anonymous):

yeah, haha

jimthompson5910 (jim_thompson5910):

like i said, it's about practice

jimthompson5910 (jim_thompson5910):

well you would use logs to isolate t

jimthompson5910 (jim_thompson5910):

sound familiar?

OpenStudy (anonymous):

OMG... all i remember is i HATED logs

jimthompson5910 (jim_thompson5910):

yeah they are a pain at first

jimthompson5910 (jim_thompson5910):

you apply the log to both sides to get 1.39377 = (1.021)^(2*t) log( 1.39377 ) = log( (1.021)^(2*t) )

jimthompson5910 (jim_thompson5910):

that allows you to pull down the exponent like so log( 1.39377 ) = 2t*log( 1.021 )

OpenStudy (anonymous):

on the left: .14419111 roughly

jimthompson5910 (jim_thompson5910):

then you divide both sides by 2*log(1.021) to get log( 1.39377 ) = 2t*log( 1.021 ) log( 1.39377 )/(2*log( 1.021 )) = t t = log( 1.39377 )/(2*log( 1.021 )) t = ??

jimthompson5910 (jim_thompson5910):

yeah you could evaluate each log one at a time like that

OpenStudy (anonymous):

how would you figure out the right side?

jimthompson5910 (jim_thompson5910):

divide both sides by 2*log(1.021)

jimthompson5910 (jim_thompson5910):

to isolate t

OpenStudy (anonymous):

how would i plug it in to a calculator?

jimthompson5910 (jim_thompson5910):

evaluate 2*log(1.021) to get _____

jimthompson5910 (jim_thompson5910):

then you divide .14419111 by that result so .14419111/(whatever you got)

OpenStudy (anonymous):

i got 2.042

OpenStudy (anonymous):

for the right side

jimthompson5910 (jim_thompson5910):

hmm something went wrong

jimthompson5910 (jim_thompson5910):

you just doubled 1.021 by mistake

jimthompson5910 (jim_thompson5910):

and forgot about the log

OpenStudy (anonymous):

i did 2*log(1.021)

jimthompson5910 (jim_thompson5910):

you left off the log though i think

OpenStudy (anonymous):

possibly, let me do it again

jimthompson5910 (jim_thompson5910):

ok

OpenStudy (anonymous):

.31005573?

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

UGHHHHH FRUSTRATION with myself

jimthompson5910 (jim_thompson5910):

its fine, as long as you can spot where you went wrong

OpenStudy (anonymous):

i did, i just put it in the calculator wrong

jimthompson5910 (jim_thompson5910):

i gotcha

OpenStudy (anonymous):

i got 7.98778888

jimthompson5910 (jim_thompson5910):

me too

jimthompson5910 (jim_thompson5910):

so about 8 years

OpenStudy (anonymous):

but there are these two choices that are very close

jimthompson5910 (jim_thompson5910):

which ones

OpenStudy (anonymous):

8 years 0 months 8 years, 1 month

jimthompson5910 (jim_thompson5910):

the most accurate I get the answer to be is 7.987787029

OpenStudy (anonymous):

it would just be 8 years, cause it just barely makes 8 years

jimthompson5910 (jim_thompson5910):

that rounds to 8 years you could go another month, but that will definitely be over $5,000

OpenStudy (anonymous):

okay thanks again man

jimthompson5910 (jim_thompson5910):

np